How many moles of NH3 can be produced from 3.76 moles of nitrogen in the following reaction: N2(g)+3H2(g)=>2NH3(g)
1
Expert's answer
2019-04-17T04:18:43-0400

n(NH3)=n(N2)*2=3.76*2=7.52 moles

(based on chemical reaction rates)